2Ω 50 1) For the circuit on the right: a) Find the current (1) going through the 45Ω resistor. b) Find Vo c) Find the power delivered by the battery. Vo 450 60V 120 30 10
Answer #1

1) current through 12 ohm = V/R = 60/12 = 5 A

current thorugh 12 ohm = 5 = I(total) * 45 / ( 45+5+12+1)

= I(total) = 7A

so I = 7 - 5= 2 A

b) current = 7A

5,12,1 ohm are in series

R1 = 5 + 12 + 1 = 18 ohm

18 and 45 ohm are in parallel

R2 = 18*45 / ( 18 + 45) = 12.857 ohm

12.857 , 2 , 3 ohm are in series

R(equi ) = 2 + 3 + 12.857 = 17.857 ohm

Vo = I*R(equi) = 7*17.857 = 125 V

c) power supplied by battery = VI = 125*7 = 875W